# TokenRec: Learning to Tokenize ID for LLM-based Generative Recommendations

### v2

@@ III-D Efficiency Evaluation
In this subsection, we analyze the inference efficiency of TokenRec compared with the general LLM-based recommendation methods (i.e., P5, CID, and POD), which generate the tokens (e.g., IDs, titles, and descriptions) for Top-$K$ items by using auto-regressive decoding and beam search. Our proposed method discards this text-decoding generation solution and introduces a generative retrieval paradigm to perform collaborative recommendations. The results, as presented in Table V, demonstrate that TokenRec can achieve superior inference efficiency with a significant improvement of approximately 1306.5% compared to the LLM-based recommendation baselines. This can be attributed to our generative retrieval paradigm, bypassing the most time-consuming auto-regressive decoding and beam search processes of LLMs [25].
TABLE V: Average inference time (milliseconds) per user.
@@ III-E Ablation Study
We can have the following observations from the ablation results in Table VI. First, each component in our approach contributes to the overall performance since eliminating any of them would result in performance degradation. Second, the masking operation and the $K$-way framework introduced in our approach not only enhance the generalizability, as demonstrated in Table IV, but also result in moderate performance improvements during the ablation experiment. Finally, the removal of the high-order collaborative knowledge learned by advanced GNNs resulted in a significant drop in performance. This indicates the importance of incorporating such knowledge for the alignment of LLMs and personalized recommendations.
TABLE VI: Results of Ablation Studies
@@ III-F2 Effect of Codebook Settings $K$ and $L$
  • With the increase of codebook depth, a progressive performance improvement of our model for recommendations is witnessed in all datasets. Nevertheless, when $K>3$, the improvement of recommendation performance becomes relatively marginal. As a result, $K=3$ is suggested by considering the trade-off between effectiveness and efficiency.
@@ IV-A Collaborative Filtering
In order to provide personalized recommendations that accord with user preferences, collaborative filtering (CF) serves as a representative technique for modeling collaborative information in recommender systems, such as user-item interactions, to capture similar user patterns and predict future interactions [41, 42]. As a representative example, matrix factorization (MF) [4, 43] vectorizes users and items into dense representations and models their interactions (i.e., user-item ID matrix) by calculating the inner products between the vectorized representations. Subsequently, NeuCF [34] incorporates neural networks with MF to decompose the use-item interactions into two low-rank matrices representing user&item embeddings. Later on, DSCF [44] takes advantage of deep language models to enhance user representations for collaborative recommendations by capturing auxiliary information from neighbors in social networks. Due to the superior representation learning capability in graphs, Graph Neural Networks (GNNs), such as LightGCN [5], GTN [6], and LTGNN [29], are proposed to capture high-order collaborative knowledge on user-item interaction graphs for enhancing the performance of recommender systems. More specifically, GNNs take advantage of the graph-structured nature of user-item interactions and model similar user behavior patterns toward items through information propagation on user-item interaction graphs. For example, GraphRec [3, 45] introduces a graph attention network-based framework to encode user-item interactions and user-user social relations for social recommendations. LightGCN [5] is introduced to largely simplify the GNN-based recommendation methods by removing feature transformation and nonlinear activation, achieving state-of-the-art prediction performance for recommendations. In addition, GTN [45] and LTGNN [29] provide an improvement by capturing the adaptive reliability and high-order linear-time patterns of interactions, respectively.
@@ IV-B LLM-based Recommender Systems
With the rapid development of Large Language Models (LLMs), such as ChatGPT and GPT-4, notable milestones have been showcased for revolutionizing natural language processing techniques [46, 47]. Technically, the unprecedented power of LLMs can be attributed to the scaling up of model/parameter size alongside a tremendous amount of training corpus [48, 49]. In particular, LLMs equipped with billion-scale parameters have exhibited unprecedented language understanding and generation ability, along with remarkable generalization capability and reasoning skills that facilitate LLMs to better generalize to unseen tasks and domains [50]. Given the emerging trends and aforementioned advancements of LLMs, LLM-empowered recommender systems have drawn increasing attention from recent studies and demonstrated distinctive abilities for advancing recommender systems [2, 25]. Notably, to harness the distinctive capabilities of LLMs for advancing recommender systems, existing studies have actively investigated various paradigms [8], including pre-training, fine-tuning, and prompting (e.g., in-context learning, chain-of-thought, and instruction tuning), for adapting LLMs to recommendation tasks. For example, P5 [10] introduces an LLM-based recommendation model that unifies diverse recommendation tasks by multi-task prompt-based pre-training, which achieves impressive zero-shot generalization capability to unseen recommendation tasks with the help of personalized prompts.
@@
Despite effectiveness, most LLM-based recommendation methods still have an intrinsic limitation on indexing users and items’ IDs in language models. The naive approach known as Independent Indexing (IID) directly assigns special tokens to represent users and items within language models. Moreover, as a natural solution, textual title indexing is proposed to utilize textual contents (e.g., titles and description) to tokenize items using LLMs’ in-vocabulary tokens [9, 11], such as the example ”Apple iPhone 15, 256 GB, black”. P5 [10] and POD [39] apply positional and whole-word embeddings to highlight the tokens representing items and users. Most recently, several studies borrow the conception of soft prompt and utilize exogenous tokens with continuous embeddings to represent users and items in LLM-based recommendations [14, 13]. Although the use of title indexing and whole-word embeddings can mitigate the issue of vocabulary explosion, it fails to capture high-order collaborative knowledge effectively and lacks generalizability for recommending to unseen users or items, as seen in P5 [10] and POD [39]. Furthermore, the discrete nature of tokens in language models presents a challenge in achieving tight alignment of LLMs in recommendations when utilizing continuous representations. Additionally, most LLM-based RecSys suffer from time-consuming inference due to auto-regressive decoding and beam search. However, the majority of existing methods for users&items tokenization fail to capture high-order collaborative knowledge in personalized recommendations effectively. Thus, this paper proposes a novel framework for LLM-based recommendations, TokenRec, which not only introduces a generalizable ID tokenization strategy to capture high-order collaborative knowledge but also proposes a generative retrieval paradigm to generate top-$K$ items efficiently.

### v3

@@ III-D Efficiency Evaluation
In this subsection, we analyze the inference efficiency of TokenRec compared with the representative LLM-based recommendation methods (i.e., P5, CID, POD, and TIGER), which generate the tokens (e.g., IDs, titles, and descriptions) for Top-$K$ items by using typical auto-regressive decoding and beam search. Our proposed method discards this text-decoding generation solution and introduces a generative retrieval paradigm to perform collaborative recommendations. This evaluation is conducted under the same default settings as those used in the overall performance comparisons. The top-20 retrieval results, as presented in Table V, demonstrate that TIGER outperforms its text-based counterparts (i.e., P5, CID, and POD) in terms of efficiency, as it utilizes compressed semantic IDs, which require fewer tokens than item titles or descriptions. More importantly, TokenRec can achieve superior inference efficiency with a significant improvement of approximately 1259.81% compared to the baselines. This can be attributed to our generative retrieval paradigm, bypassing the most time-consuming auto-regressive decoding and beam search processes of LLMs [29].
TABLE V: Comparison of the average inference time per user (in milliseconds) for Top-20 recommendations.
@@ III-E Ablation Study
  • s K-Means, VQ-VAE, and RQ-VAE: Substitute the proposed MQ-Tokenizers with three typical vector quantization/clustering methods, specifically K-Means, VQ-VAE [19], and RQ-VAE [48] to validate its effectiveness.
We can have the following observations from the ablation results in Table VI. First, each component in our approach contributes to the overall performance since eliminating any of them would result in performance degradation. Second, the masking operation and the $K$-way framework introduced in our approach not only enhance the generalizability, as demonstrated in Table IV, but also result in moderate performance improvements during the ablation experiment. Moreover, the removal of the high-order collaborative knowledge learned by lightGCN resulted in a significant drop in performance. This indicates the importance of incorporating such knowledge for the alignment of LLMs with personalized recommendations. Furthermore, it demonstrates that the quality of collaborative embeddings is crucial to the effectiveness of our tokenizer and model. Third, the comparison with the two representative quantization methods, namely VQ-VAE and RQ-VAE, illustrates the effectiveness of our MQ-Tokenizers in encoding collaborative knowledge for LLM-based recommendations, while the inferior performance of K-Means underscores the necessity for a more robust quantization approach.
TABLE VI: Results of Ablation Studies.
@@ III-F2 Effect of Codebook Settings $K$ and $L$
  • With the increase of codebook depth, a progressive performance improvement of our model for recommendations is witnessed in all datasets. Nevertheless, when $K>3$, the improvement of recommendation performance becomes relatively marginal. Thus, $K=3$ is suggested by considering the trade-off between effectiveness and efficiency.
@@ IV-A Collaborative Filtering
In order to provide personalized recommendations that accord with user preferences, collaborative filtering (CF) serves as a representative technique for modeling collaborative information in recommender systems, such as user-item interactions, to capture similar user patterns and predict future interactions [49, 50, 51]. As a representative example, matrix factorization (MF) [5, 52] vectorizes users and items into dense representations and models their interactions (i.e., user-item ID matrix) by calculating the inner products between the vectorized representations. Subsequently, NeuCF [39] incorporates neural networks with MF to decompose the use-item interactions into two low-rank matrices representing user&item embeddings. Later on, DSCF [53] takes advantage of deep language models to enhance user representations for collaborative recommendations by capturing auxiliary information from neighbors in social networks. Due to the superior representation learning capability in graphs, Graph Neural Networks (GNNs), such as LightGCN [6], GTN [7], and LTGNN [34], are proposed to capture high-order collaborative knowledge on user-item interaction graphs for enhancing the performance of recommender systems. More specifically, GNNs take advantage of the graph-structured nature of user-item interactions and model similar user behavior patterns toward items through information propagation on user-item interaction graphs. For example, GraphRec [4, 54] introduces a graph attention network-based framework to encode user-item interactions and user-user social relations for social recommendations. LightGCN [6] is introduced to largely simplify the GNN-based recommendation methods by removing feature transformation and nonlinear activation, achieving state-of-the-art prediction performance for recommendations. In addition, GTN [54] and LTGNN [34] provide an improvement by capturing the adaptive reliability and high-order linear-time patterns of interactions, respectively.
@@ IV-B LLM-based Recommender Systems
With the rapid development of Large Language Models (LLMs), such as ChatGPT and GPT-4, notable milestones have been showcased for revolutionizing natural language processing techniques [55, 56, 57, 58]. In particular, LLMs equipped with billion-scale parameters have exhibited unprecedented language understanding and generation ability, along with remarkable generalization capability and reasoning skills that facilitate LLMs to better generalize to unseen tasks and domains [59]. Given the emerging trends and aforementioned advancements of LLMs, LLM-empowered recommender systems have drawn increasing attention from recent studies and demonstrated distinctive abilities for advancing recommender systems [3, 29, 60, 61]. Notably, to harness the distinctive capabilities of LLMs for advancing recommender systems, existing studies have actively investigated various paradigms [9], including pre-training, fine-tuning, and prompting, for adapting LLMs to recommendation tasks. For example, P5 [12] introduces an LLM-based recommendation model that unifies diverse recommendation tasks by multi-task prompt-based pre-training, which achieves impressive zero-shot generalization capability to unseen recommendation tasks with personalized prompts.
@@
Despite their effectiveness, most LLM-based recommendation methods still have an intrinsic limitation on indexing users and items’ IDs in language models. The straightforward strategy, Independent Indexing (IID), involves directly allocating unique tokens to users and items. However, it proves impractical and unfeasible for large-scale real-world recommendation systems, where the user and item populations can easily extend into the billions, significantly inflating the token vocabulary within LLMs. As a natural solution, textual title indexing is proposed to utilize textual contents (e.g., titles and description) to tokenize items using LLMs’ in-vocabulary tokens [11, 13], such as the example “Apple iPhone 15, 256 GB, black”. Moreover, P5 [12] and POD [45] apply positional and whole-word embeddings to highlight the tokens representing items and users. More recently, TIGER [46] effectively condenses copious textual data into a few semantic IDs using a residual vector quantization method. By utilizing these quantized semantic IDs as tokens, TIGER is constructed as a Sequence-to-Sequence Transformer-based model for sequential recommendation tasks. Although the uses of title indexing, whole-word embeddings, and semantic IDs can mitigate the issue of vocabulary explosion, they fail to capture high-order collaborative knowledge effectively and lack generalizability for recommending to unseen users or items. To incorporate such valuable knowledge, several studies, including CoLLM [16], LlaRA [15], and E4SRec [62], borrow the conception of soft prompt and utilize exogenous tokens with continuous embeddings to represent users and items in LLM-based recommendations. However, the discrete nature of tokens in language models presents a challenge in achieving tight alignment of LLMs in recommendations when utilizing continuous representations. Although META ID [63] has suggested integrating collaborative knowledge into discrete tokens through clustering items’ and users’ representations derived from skip-gram models, a robust tokenizer remains unspecific to ensure the effectiveness of the quantization (i.e., clustering) procedure. Additionally, most LLM-based RecSys suffer from time-consuming inference due to auto-regressive decoding and beam search. To address these under-explored issues, this paper proposes a novel framework for LLM-based recommendations, TokenRec, which not only introduces a generalizable ID tokenization strategy to capture high-order collaborative knowledge but also proposes a generative retrieval paradigm to generate top-$K$ items efficiently. Unlike existing ID indexing methods that also utilize vector quantization, such as TIGER, our tokenizer features a parallel K-way structure and incorporates a masking operation for robust ID tokenization. Moreover, it is designed to capture insightful collaborative knowledge enriched with user-item interaction information.
@@ V Conclusion
Acknowledgments
The research described in this paper has been partially supported by the National Natural Science Foundation of China (project no. 62102335), General Research Funds from the Hong Kong Research Grants Council (project no. PolyU 15207322, 15200023, 15206024, and 15224524), internal research funds from Hong Kong Polytechnic University (project no. P0042693, P0048625, and P0051361). This work was supported by computational resources provided by The Centre for Large AI Models (CLAIM) of The Hong Kong Polytechnic University.